# DAO: A New and Public Non-Relativistic Reflection Model

### v1

@@ Abstract
We present a new non-relativistic reflection model, DAO, designed to calculate reflection spectra in the rest frame of accretion disks in X-ray binaries and active galactic nuclei. The model couples the XSTAR code, which treats atomic processes, with the Feautrier method for solving the radiative transfer equation. A key feature of DAO is the incorporation of a high-temperature corrected cross section and an exact redistribution function to accurately treat Compton scattering. Furthermore, the model accommodates arbitrary illuminating spectra, enabling applications across diverse physical conditions. We investigate the spectral dependence on key physical parameters and benchmark the results against the widely used reflionx and xillver codes.
@@ 1 Introduction
The X-ray spectra of X-ray binary systems (XRBs) and active galactic nuclei (AGNs) are essential for understanding the process of accretion onto black holes (BHs). These spectra typically consist of several components: (1) thermal emission from a geometrically thin and optically thick accretion disk (Shakura1973; 1973blho.conf..343N); (2) non-thermal emission from a hot corona ($\sim$100 keV), whose spectrum follows a power-law–like shape with a high-energy cutoff (Thorne1975ApJ...195L.101T; Shapiro1976ApJ...204..187S; Zdziarski1995ApJ...438L..63Z); and (3) reprocessed coronal emission from the accretion disk, commonly referred to as the reflection component (1995Natur.375..659T; Risaliti2013Natur.494..449R; Bambi2021SSRv..217...65B). The reflection spectrum encodes information about the disk properties (e.g., density, ionization state, and elemental abundances) as well as the spacetime geometry near the central black hole (1989MNRAS.238..729F; George1991MNRAS.249..352G; 2001MNRAS.327...10B; 2010ApJ...718..695G; 2010MNRAS.409.1534D; 2019ApJ...874..135T; 2019ApJ...875...56T; 2021ApJ...913...79T). Modeling the reflection spectrum and fitting it to the data is particularly useful for extracting this information (Bambi2021SSRv..217...65B; Reynolds2021ARA&A..59..117R; 2023ApJ...951..145L; Liu2023ApJ...950....5L; Draghis2024ApJ...969...40D).
@@
Calculating the disk reflection spectrum involves two steps: (1) computing the reflection spectrum in the local rest frame by solving the radiative transfer equation, taking into account processes such as photoionization and Compton scattering (George1991MNRAS.249..352G; 2010ApJ...718..695G); and (2) integrating the local reflection spectrum from the inner to the outer disk while including all relativistic effects, such as light bending, Doppler shifts, and gravitational redshift (1989MNRAS.238..729F; 2010MNRAS.409.1534D; 2025MNRAS.536.2594L; 2025ApJ...989..168H). Integration can be done based on a standard pre-tabulated transfer function (see 2024arXiv240812262B, for more details).
@@
For the rest-frame reflection spectrum, reflionx [^1] https://github.com/honghui-liu/reflionx_tables (1978ApJ...219..292R; 1993MNRAS.261...74R; 2001MNRAS.327...10B; 2005MNRAS.358..211R; 2007MNRAS.381.1697R) and xillver [^2] https://www.sternwarte.uni-erlangen.de/~dauser/research/relxill/ (2010ApJ...718..695G; 2011ApJ...731..131G; 2013ApJ...768..146G) are the most widely used models in X-ray reflection data analysis. Another code, TITAN (2003A&A...407...13D), is also used in the literature. The main difference between them lies in the treatment of atomic physics and radiative transfer. xillver adopts the Feautrier method for solving the radiative transfer equation, allowing the calculation of emergent spectra at different viewing angles. It also includes fluorescent lines that are missing in reflionx and TITAN, due to the powerful atomic database of XSTAR (2001ApJS..133..221K). In addition, TITAN only includes Compton heating and cooling in the energy balance, without accounting for Comptonization in the redistribution of photon energy (2000A&A...357..823D). Both reflionx and xillver treat Compton scattering using the same Gaussian-approximated redistribution function, which is only valid to describe the energy redistribution for low-energy photons ($E_{\rm ph}\ll m_{e}c^{2}$) scattered by low-temperature electron clouds ($kT_{e}\ll E_{\rm ph}$).
@@
Energy redistribution in the Compton scattering process should be described by the full quantum-mechanical treatment, and the electrons in hot plasma should follow a relativistic Maxwellian distribution (1993AstL...19..262N). For scattering of low-energy photons by a low-temperature electron population, both the Gaussian and exact redistribution functions maintain nearly symmetric profiles. In other cases, the energy-exchange probabilities of photons differ markedly between the two formulations (see Appendix B). 2020ApJ...897...67G investigated this effect on the reflection spectrum, but it has not yet been implemented in the xillver table model. They compared the spectrum calculated with the exact Compton scattering kernel with the publicly available xillver table and the Monte Carlo model pexrav (1995MNRAS.273..837M). The results demonstrate that the commonly used approximations for Compton scattering can lead to significant deviations in the high-energy band.
@@
In addition, 2018A&A...614A..79P reported a noticeable discrepancy above 30 keV when comparing xillverCp with pexrav/pexriv. They attributed this difference to the fact that xillverCp does not account for the Klein–Nishina effect, which effectively absorbs part of the energy of the incident photons at $E\gtrsim 50$ keV (2003MNRAS.342..355Z).
@@
In this paper, we present a new non-relativistic reflection model, DAO. The model employs the latest version of the photoionization code XSTAR v2.59, to compute atomic processes, thermal equilibrium, and ionization balance, and incorporates the Feautrier method to solve the radiative transfer. It includes the exact redistribution function for Compton scattering as explained in 2020ApJ...897...67G. We adopt a constant-density gas with a specified column density under a plane-parallel geometry and azimuthal symmetry. The model is open source and designed to be highly flexible, allowing users to adjust nearly all parameters. DAO is available on GitHub [^3] https://github.com/ABHModels/DAO and Zenodo (huang_2025_17839597).
@@
The manuscript is organized as follows. In Section 2, we introduce the methodology of our model (radiative transfer, iteration process, etc.). The default configuration, as detailed in Section 2.1, is adopted throughout this work unless otherwise specified. In Section 3, we present a comparative analysis of our model against reflionx and xillver. We also investigate the impact of various parameters, such as the ionization parameter, incident spectrum, incidence angle, and hydrogen density, on the final reflection spectrum. Finally, Section 4 summarizes our findings and outlines future research directions.
2 METHOD
@@
2.1 Basic setup
We consider a constant density slab illuminated from above and below. The slab represents the disk atmosphere, and the illumination from above represents emission from the corona irradiating the disk atmosphere, at an incidence angle $\theta_{i}$. DAO provides several standard forms for the incident spectrum, including a power law, a power law with a high-energy cutoff, the thermal Comptonisation model nthComp (1996MNRAS.283..193Z), and a blackbody. For the exponentially cut off power law, the specific intensity is given by [^4] Note that our standard power law form does not include the 0.1 keV cut off that is adopted in xillver.
@@ 2.1 Basic setup
and the nthcomp spectrum is parameterised by $\Gamma$, the seed photon temperature $kT_{seed}$ (hardwired to $0.1$ keV for a multi-temeprature seed photon spectrum), and the electron temperature $kT_{E}$. These spectral forms have been widely used in the analysis of data from accreting black holes and neutron stars. Physical processes such as returning radiation and Comptonization can further modify the shape of the illumination (1999MNRAS.309..561Z; 2021ApJ...910...49R; 2022MNRAS.514.3965D; 2024ApJ...965...66M; 2024ApJ...976..229M). To enable broader exploration of physical scenarios, DAO also allows user-defined incident spectra. The incident spectrum at the upper boundary is normalised by the ionization parameter (1969ApJ...156..943T)
@@
where $F_{x}$ is the illuminating X-ray flux [^5] Our model calculates $F_{x}$ in the range of 0.1-1000 keV, such as xillver (2022ApJ...926...13G) but different from reflionx, which defines $F_{x}$ from 0.0013 to 1000 keV and $n_{h}$ is the hydrogen number density.
DAO includes the option to incorporate a blackbody radiation source at the lower boundary, corresponding to the illumination of the disk atmosphere by thermal emission from the disk midplane. This feature is crucial for XRBs, as their accretion disks can be sufficiently hot to produce strong ionizing radiation (Liu2023ApJ...950....5L). A similar boundary condition is implemented in reflionx (2007MNRAS.381.1697R). At present, such bottom illumination is currently not included in the xillver table model. Therefore, to facilitate a direct comparison, we set the lower boundary condition to zero in this work.
@@
The grid-related parameters define the discretization of energy, angle, and optical depth, as well as the maximum number of iteration steps. In this work, we adopt 10 linearly spaced cosine-angle bins ranging from $0.05$ to $0.95$ for the angular quadrature. For the remaining dimensions, a Thomson optical-depth grid consisting of 200 points logarithmically spaced from $10^{-4}$ to $10$ is used, together with a default energy grid of 5000 points covering the range from $0.1\,\mathrm{eV}$ to $1000\,\mathrm{keV}$. The maximum number of main iteration steps, $N_{\mathrm{main}}$, and radiative-transfer iteration steps are set to 15 and 200, respectively. These values are sufficient to ensure convergence of the model, as discussed in detail in Section 2.4.
@@
2.2 Radiative transfer equation
This section outlines the methodology employed in our reflection model DAO. We employ XSTAR v2.59 (2021ApJ...908...94K) to treat the relevant atomic processes, and adopt the Feautrier method (hereafter FTM) (1978stat.book.....M; 2015tsaa.book.....H) to iteratively solve the radiative transfer equation.
@@ 2.2 Radiative transfer equation
FTM is based on the solution of the second-order radiative transfer equation (1905ApJ....21....1S), which is derived by considering two rays propagating in opposite directions, denoted by $\pm\mu$. From these, we define the mean-like intensity at depth $z$ and in the range $0\leq\mu\leq 1$ as
@@
A key assumption in this equation is the forward–backward symmetry of the source function [^7] In a real disk irradiated by a compact corona, the symmetry between forward and backward directions is broken because the radiation is strongly anisotropic., i.e., $S(+\mu)=S(-\mu)$. This symmetry is essential for solving the second-order radiative transfer equation (Eq. 7) using the FTM. When this approximation is not valid, alternative methods must be employed (1975ApJ...202..250M). The source function can then be expressed as
@@
The second term on the right-hand side in Eq. 8 represents the Compton scattering of photons by electrons in the gas. Unlike the Klein–Nishina scattering cross section used in previous reflection models (2005MNRAS.358..211R), which is only valid for low-temperature electrons (see Figure 1 in 2020ApJ...897...67G), our approach accounts for high-temperature gas by employing the Compton scattering cross section averaged over a relativistic Maxwellian electron distribution (1996ApJ...470..249P). The mathematical form of $\sigma_{cs}$ can be found in Eq. B10 in Appendix B. $J^{c}$ is the mean intensity after scattering:
@@
where $E_{i}$ and $E_{f}$ are the photon energies before and after scattering, respectively, and $J$ is the mean intensity before scattering. The Compton scattering redistribution function, $R$, satisfies the detailed balance condition relating the initial and final photon energies, $E_{i}$ and $E_{f}$, respectively. This condition holds for a Maxwellian electron distribution and is consistent with the photons following a Wien distribution in thermal equilibrium (see 1973erh..book.....P, Chapter VIII):
@@
Note that the photon phase-space factors are implicitly absorbed into the definition of $R$. We adopt the normalization condition given by (1996ApJ...470..249P; 2017MNRAS.469.2032M):
@@
The redistribution function $R$ governs the change in photon energy during the scattering process (see Appendix B). In addition to altering photon energies, Compton scattering also modifies photon propagation directions. At the present stage, we employ the angle-averaged redistribution function, consistent with the approaches used in xillver and reflionx. 2012A&A...545A.120S showed that there exists a $\sim 2\%$ difference between the angle-dependent and angle-averaged redistribution functions in a pure hydrogen atmosphere (see their Figure 2). In future work, we plan to incorporate the full angular dependence of directional changes.
Previous public reflection models have employed a Gaussian-approximated redistribution function. In this work, we compare DAO with other models using both the Gaussian approximation and a more rigorous second-order exact redistribution function (see Section 3.1). We further quantify the deviations introduced by the Gaussian approximation in Section 3.2. The mathematical forms of both the second-order exact and Gaussian-approximated redistribution functions are provided in Appendix B. To facilitate flexibility, DAO includes a built-in selector that allows users to switch between the two redistribution schemes.
2.3 Boundary condition
The second-order radiative transfer equation (Eq. 7) can be reformulated as a block tridiagonal system of the form $T\cdot j=R$. The detailed derivation of this transformation is provided in Appendix A. To solve this system efficiently, we employ the Thomas algorithm, which performs recursive forward elimination followed by back substitution.
The disk surface is irradiated by hard X-ray photons from the corona, while the disk itself emits thermal radiation. This configuration can be formulated as a two-boundary problem, making it suitable for solving the radiative transfer equation (Eq. 7). The upper boundary condition is determined by the incident radiation from an external source such as the corona. Following 1967ApJ...150L..53A, we adopt the second-order exact boundary condition derived from a Taylor series (see Appendix A):

### v2

@@ Abstract
We present a new non-relativistic reflection model, DAO, designed to calculate reflection spectra in the rest frame of accretion disks in X-ray binaries and active galactic nuclei. Employing the mathematical formalism of the xillver model, DAO couples the XSTAR code, which treats atomic processes, with the Feautrier method for solving the radiative transfer equation. A key feature of DAO is that it is open source and allows users to specify arbitrary illuminating spectra, enabling applications across diverse physical conditions. We incorporate a high-temperature corrected cross section and an exact redistribution function to accurately treat Compton scattering, and we use the most complete and up-to-date atomic database available. We investigate the spectral dependence on key physical parameters and benchmark the results against the publicly available table models for the widely used reflionx and xillver codes.
@@ I Introduction
The X-ray spectra of X-ray binary systems (XRBs) and active galactic nuclei (AGNs) are essential for understanding the process of accretion onto black holes (BHs). These spectra typically consist of several components: (1) thermal emission from a geometrically thin and optically thick accretion disk (Shakura and Sunyaev 1973; Novikov and Thorne 1973); (2) non-thermal emission from a hot corona ($\sim$100 keV), whose spectrum follows a power-law–like shape with a high-energy cutoff (Thorne and Price 1975; Shapiro et al. 1976; Zdziarski et al. 1995); and (3) reprocessed coronal emission from the accretion disk, commonly referred to as the reflection component (Tanaka et al. 1995; Risaliti et al. 2013; Bambi et al. 2021). The reflection spectrum encodes information about the disk properties (e.g., density, ionization state, and elemental abundances) as well as the spacetime geometry near the central black hole (Fabian et al. 1989; George and Fabian 1991; Ballantyne et al. 2001; García and Kallman 2010; Dauser et al. 2010; Tripathi et al. 2019b, a, 2021). Modeling the reflection spectrum and fitting it to the data is particularly useful for extracting this information (Bambi et al. 2021; Reynolds 2021; Liu et al. 2023b, a; Draghis et al. 2024).
@@
Calculating the disk reflection spectrum involves two steps: (1) computing the reflection spectrum in the local rest frame by solving the radiative transfer equation, taking into account processes such as photoionization and Compton scattering (George and Fabian 1991; García and Kallman 2010); and (2) integrating the local reflection spectrum from the inner to the outer disk while including all relativistic effects, such as light bending, Doppler shifts, and gravitational redshift (Fabian et al. 1989; Dauser et al. 2010; Liu et al. 2025; Huang et al. 2025b). Integration can be done based on a standard pre-tabulated transfer function (see Bambi 2024, for more details).
@@
For the rest-frame reflection spectrum, reflionx [^1] https://github.com/honghui-liu/reflionx_tables (Ross et al. 1978; Ross and Fabian 1993; Ballantyne et al. 2001; Ross and Fabian 2005, 2007) and xillver [^2] https://www.sternwarte.uni-erlangen.de/~dauser/research/relxill/ (García and Kallman 2010; García et al. 2011, 2013) are the most widely used models in X-ray reflection data analysis. Another code, TITAN (Dumont et al. 2003), is also used in the literature. The main difference between them lies in the treatment of atomic physics and radiative transfer. xillver adopts the Feautrier method for solving the radiative transfer equation, allowing the calculation of emergent spectra at different viewing angles. It also includes fluorescent lines that are missing in reflionx and TITAN, due to the powerful atomic database of XSTAR (Kallman and Bautista 2001). In addition, TITAN only includes Compton heating and cooling in the energy balance, without accounting for Comptonization in the redistribution of photon energy (Dumont et al. 2000). The currently available reflionx and xillver table models employ distinct approaches to Compton scattering. In the reflionx model, Ross et al. (1978); Ross and Fabian (1993, 2005) utilize the Fokker-Planck equation (Kompaneets 1957) to address multiply scattered photons, while applying escape probabilities for unscattered and once-scattered line photons. For once-scattered line photons, Ross et al. (1978) approximate the energy redistribution using a Gaussian redistribution function. In the xillver model, García and Kallman (2010); García et al. (2013, 2014) employ a Gaussian redistribution function to account for the energy shift in Compton scattering for all photons. Although Nayakshin et al. (2000) demonstrated the validity of treating all photons via the Gaussian approximation, this approximation breaks down at high photon energies ($E_{\rm ph}\gg m_{e}c^{2}$) or high electron temperatures ($kT_{e}\gg E_{\rm ph}$).
@@
Energy redistribution in the Compton scattering process should be described by the full quantum-mechanical treatment, and the electrons in hot plasma should follow a relativistic Maxwellian distribution (Nagirner and Poutanen 1993). For scattering of low-energy photons by a low-temperature electron population, both the Gaussian and exact redistribution functions maintain nearly symmetric profiles. In other cases, the energy-exchange probabilities of photons differ markedly between the two formulations (see Appendix B). García et al. (2020) investigated this effect on the reflection spectrum, and implemented it onto the xillver code, but table models including this new feature have not yet been publicly released and the xillver source code is not currently publicly available. They compared the spectrum calculated with the exact Compton scattering kernel with the publicly available xillver table and the Monte Carlo model pexrav (Magdziarz and Zdziarski 1995). The results demonstrate that the commonly used approximations for Compton scattering can lead to significant deviations in the high-energy band.
@@
In addition, Poutanen et al. (2018) reported a noticeable discrepancy above 30 keV when comparing xillverCp with pexrav/pexriv. They attributed this difference to the fact that xillverCp does not account for the Klein–Nishina effect, which effectively absorbs part of the energy of the incident photons at $E\gtrsim 50$ keV (Zdziarski et al. 2003).
@@
In this paper, we present a new non-relativistic reflection code, DAO. We employ the xillver implementation: employing the photoionization code XSTAR to compute atomic processes, thermal equilibrium, and ionization balance, and incorporating the Feautrier method to solve the radiative transfer. We include the exact redistribution function for Compton scattering provided by García et al. (2020). We adopt a constant-density gas with a specified column density under a plane-parallel geometry and azimuthal symmetry. A key benefit of DAO is that it is open source and designed to be highly flexible, allowing users to adjust nearly all parameters. This will enable any user to, for example, create model grids for any arbitrary illuminating spectrum, or to implement more advanced model assumptions. A secondary benefit is the insight that can be gained from comparing the outputs of DAO and xillver. Since the mathematical formalism is identical, the two models should in principle agree exactly, although in practice we use different versions of ingredients such as XSTAR. DAO is available on GitHub [^3] https://github.com/ABHModels/DAO and Zenodo (Huang et al. 2025a).
@@
The manuscript is organized as follows. In Section II, we introduce the methodology of our model (radiative transfer, iteration process, etc.). The default configuration, as detailed in Section II.1, is adopted throughout this work unless otherwise specified. In Section III, we present a comparative analysis of our model against reflionx and xillver. We also investigate the impact of various parameters, such as the ionization parameter, incident spectrum, incidence angle, and hydrogen density, on the final reflection spectrum. Finally, Section IV summarizes our findings and outlines future research directions.
II METHOD
@@
II.1 Basic setup
We consider a constant density slab illuminated from above and below. The slab represents the disk atmosphere, and the illumination from above represents emission from the corona irradiating the disk atmosphere, at an incidence angle $\theta_{i}$. DAO provides several standard forms for the incident spectrum, including a power law, a power law with a high-energy cutoff, the thermal Comptonisation model nthComp (Zdziarski et al. 1996), and a blackbody. For the exponentially cut off power law, the specific intensity is given by [^4] Note that our standard power law form does not include the 0.1 keV cut off that is adopted in xillver.
@@ II.1 Basic setup
and the nthcomp spectrum is parameterised by $\Gamma$, the seed photon temperature $kT_{seed}$ (hardwired to $0.1$ keV for a multi-temeprature seed photon spectrum), and the electron temperature $kT_{E}$. These spectral forms have been widely used in the analysis of data from accreting black holes and neutron stars. Physical processes such as returning radiation and Comptonization can further modify the shape of the illumination (Życki et al. 1999; Riaz et al. 2021; Dauser et al. 2022; Mirzaev et al. 2024b, a). To enable broader exploration of physical scenarios, DAO also allows user-defined incident spectra. The incident spectrum at the upper boundary is normalised by the ionization parameter (Tarter et al. 1969)
@@
where $F_{x}$ is the illuminating X-ray flux [^5] Our model calculates $F_{x}$ in the range of 0.1-1000 keV, such as xillver (García et al. 2022) but different from reflionx, which defines $F_{x}$ from 0.0013 to 1000 keV and $n_{h}$ is the hydrogen number density.
DAO includes the option to incorporate a blackbody radiation source at the lower boundary, corresponding to the illumination of the disk atmosphere by thermal emission from the disk midplane. This feature is crucial for XRBs, as their accretion disks can be sufficiently hot to produce strong ionizing radiation (Liu et al. 2023a). A similar boundary condition is implemented in reflionx (Ross and Fabian 2007). At present, such bottom illumination is currently not included in the xillver table model. Therefore, to facilitate a direct comparison, we set the lower boundary condition to zero in this work.
@@
The grid-related parameters define the discretization of energy, angle, and optical depth, as well as the maximum number of iteration steps. In this work, we adopt 10 linearly spaced cosine-angle bins ranging from $0.05$ to $0.95$ for the angular quadrature. For the remaining dimensions, a Thomson optical-depth grid consisting of 200 points logarithmically spaced from $10^{-4}$ to $10$ is used, together with a default energy grid of 5000 points covering the range from $0.1\,\mathrm{eV}$ to $1000\,\mathrm{keV}$. The maximum number of main iteration steps, $N_{\mathrm{main}}$, and radiative-transfer iteration steps are set to 15 and 200, respectively. These values are sufficient to ensure convergence of the model, as discussed in detail in Section II.4.
@@
II.2 Radiative transfer equation
This section outlines the methodology employed in our reflection model DAO, which follows the mathematical formalism employed for the xillver models (García and Kallman 2010; García et al. 2013). We employ XSTAR v2.59 (Kallman et al. 2021) to treat the relevant atomic processes, and adopt the Feautrier method (hereafter FTM) (Mihalas 1978; Hubeny and Mihalas 2015) to iteratively solve the radiative transfer equation.
@@ II.2 Radiative transfer equation
FTM is based on the solution of the second-order radiative transfer equation (Schuster 1905), which is derived by considering two rays propagating in opposite directions, denoted by $\pm\mu$. From these, we define the mean-like intensity at depth $z$ and in the range $0\leq\mu\leq 1$ as
@@
A key assumption in this equation is the forward–backward symmetry of the source function [^7] In a real disk irradiated by a compact corona, the symmetry between forward and backward directions is broken because the radiation is strongly anisotropic., i.e., $S(+\mu)=S(-\mu)$. This symmetry is essential for solving the second-order radiative transfer equation (Eq. 7) using the FTM. When this approximation is not valid, alternative methods must be employed (Milkey et al. 1975). The source function can then be expressed as
@@
The second term on the right-hand side in Eq. 8 represents the Compton scattering of photons by electrons in the gas. Unlike the Klein–Nishina scattering cross section used in previous reflection models (Ross and Fabian 2005), which is only valid for low-temperature electrons (see Figure 1 in García et al. 2020), our approach accounts for high-temperature gas by employing the Compton scattering cross section averaged over a relativistic Maxwellian electron distribution (Poutanen and Svensson 1996). The mathematical form of $\sigma_{cs}$ can be found in Eq. B10 in Appendix B. $J^{c}$ is the mean intensity after scattering:
@@
where $E_{i}$ and $E_{f}$ are the photon energies before and after scattering, respectively, and $J$ is the mean intensity before scattering. The Compton scattering redistribution function, $R$, satisfies the detailed balance condition relating the initial and final photon energies, $E_{i}$ and $E_{f}$, respectively. This condition holds for a Maxwellian electron distribution and is consistent with the photons following a Wien distribution in thermal equilibrium (see Pomraning 1973, Chapter VIII):
@@
Note that the photon phase-space factors are implicitly absorbed into the definition of $R$. We adopt the normalization condition given by (Poutanen and Svensson 1996; Madej et al. 2017):
@@
The redistribution function $R$ governs the change in photon energy during the scattering process (see Appendix B). In addition to altering photon energies, Compton scattering also modifies photon propagation directions. At the present stage, we employ the angle-averaged redistribution function, consistent with the approaches used in xillver and reflionx. Suleimanov et al. (2012) showed that there exists a $\sim 2\%$ difference between the angle-dependent and angle-averaged redistribution functions in a pure hydrogen atmosphere (see their Figure 2). In future work, we plan to incorporate the full angular dependence of directional changes.
Previous public reflection models have employed a Gaussian-approximated redistribution function. In this work, we compare DAO with other models using both the Gaussian approximation and a more rigorous second-order exact redistribution function previously explored by García et al. (2020) (see Section III.1). We further quantify the deviations introduced by the Gaussian approximation in Section III.2. The mathematical forms of both the second-order exact and Gaussian-approximated redistribution functions are provided in Appendix B. To facilitate flexibility, DAO includes a built-in selector that allows users to switch between the two redistribution schemes.
II.3 Boundary condition
To solve the second-order radiative transfer equation (Eq. 7), we again use the same mathematical formalism used by xillver, namely FTM. Specifically, we reformulate the radiative transfer equation as a block tridiagonal system of the form $T\cdot j=R$ (see Appendix A for a detailed derivation) and then solve with the Thomas algorithm, which performs recursive forward elimination followed by back substitution.
The disk surface is irradiated by hard X-ray photons from the corona, while the disk itself emits thermal radiation. This configuration can be formulated as a two-boundary problem, making it suitable for solving the radiative transfer equation (Eq. 7). The upper boundary condition is determined by the incident radiation from an external source such as the corona. Following Auer (1967), we adopt the second-order exact boundary condition derived from a Taylor series (see Appendix A):